OpenStudy (anonymous):

If y=6 when x=10, find x when y=18 please i need all the help i can get

OpenStudy (tkhunny):

You'll have to provide more information. Some osrt of relationship must be established.

OpenStudy (anonymous):

um, supose y varies as x. Write as direct variation equation then solve?

OpenStudy (tkhunny):

"Direct Variation" s a sufficient hint. "y varies directly with x" MEANS y = kx for some value of k. Now we need an example. 6 = k*10 ==> k = 6/10 = 3/5 Now we know that y = (3/5)x and we can answer the question.